OpenStudy (anonymous):

Let f(x) = 2x - 6. Solve f-1(x) when x = 2.

OpenStudy (nurali):

To find the inverse of f(x) = 2x - 6, then let ƒ(x) be y, i.e. y = 2x - 6 Make x the subject of formula by adding 6 to both sides and then dividing by 2, which you'll have (y + 6)/2 = x the x & y variables to get the inverse of y, then (x + 6)/2 = y Remember that f(x) is y, so also the inverse of y is f-1(x), then f-1(x) = (x + 6)/2, where x = 2, you'll have f-1(2) = (2 + 6)/2 f-1(2) = 8/2 f-1(2) = 4

OpenStudy (anonymous):

Thankyou.

OpenStudy (nurali):

My Pleasure.